# Simulation of charged nanotubes self-assembly during evaporation of a sessile droplet on a substrate

### v1

@@ II.2.4 Electrostatic interaction and viscous friction
Within the scope of this proposed model, the motion of a particle in the flow is superimposed onto the motion of the particle relative to the liquid under the influence of both electrostatic and viscous forces. Consider separately the possible types of motion of the nanotube under the influence of electrostatic forces: (EI) translational motion, (EII) free rotational motion relative to the center of mass of the particle, (EIII) and rotational motion relative to a point of contact with another particle or the fixing radius, $R_{f}$. Assume that the current particle is affected only by neighboring particles located within a cutoff radius $R_{\mathrm{cut}}=l$.
Suppose that at each small time step, the nanotube under consideration undergoes translational motion (EI) with a constant velocity $\vec{v}^{rod}$ (steady motion is the result of the drag force balancing the electrostatic force),
$\vec{F}-k_{f}\vec{v}^{rod}=0,$
$\Rightarrow\quad\left.v^{rod}_{x}\right|_{i}=\left.\left(\sum\limits_{j}F_{x}^{j}\right)/k_{f}\right|_{i},\qquad\left.v^{rod}_{y}\right|_{i}=\left.\left(\sum\limits_{j}F_{y}^{j}\right)/k_{f}\right|_{i}.$ (7)
Here, $\vec{F}^{j}$ is the electrostatic force acting on the current nanotube from the $j$-th particle. The forces are summed across all neighboring nanotubes with which the current particle interacts. The viscous friction force is considered to be proportional to the first power of the particle velocity relative to the medium. The viscous friction coefficient $k_{f}$ is described in detail in Appendix B. To simplify calculations, we assume that the charge of a nanotube is concentrated at its ends (two identical point charges, $q$). The exact formula for calculating the electrostatic force in the case when the charge is evenly distributed over the nanotube can be found in Supplementary Material A. But in practice, it is difficult to use such a formula from a computational point of view.
@@
The force exerted by nanotube $I$ ($A_{1}\,A_{2}$) on nanotube $II$ ($B_{1}\,B_{2}$), is equal to the sum of four forces, each of which is determined by Coulomb’s law
@@
For example, $\vec{F}_{11}=kq^{2}\frac{\overrightarrow{A_{1}B_{1}}}{|A_{1}B_{1}|^{3}}$, where $k$ is the constant in Coulomb’s law. The rest of the forces are recorded in the same way. During the time interval $\Delta t$, the particle moves by $\left.v^{rod}_{x}\right|_{i}\cdot\Delta t$ along the $OX$ axis and by $\left.v^{rod}_{y}\right|_{i}\cdot\Delta t$ along the $OY$ axis.
Consider the free rotational motion relative to the center of mass of the nanotube (EII). Let’s write the moment of forces $\vec{M}$, acting on particle $II$ ($B_{1}\,B_{2}$) from the neighboring particle $I$ ($A_{1}\,A_{2}$), relative to the center of the nanotube corresponding to point $C_{2}$ (see Fig. 2),
@@
Let us assume that at each small time step $i$ the moment of resistance balances the moment of electrostatic force, $\sum_{j}\vec{M}^{j}=\vec{M}_{\mathrm{vis}}|_{i}$. Then, the particle $B_{1}\,B_{2}$ rotates around the center of mass at a constant angular velocity at each time step,
$\left.\omega_{z}\right|_{i}=4\left.\frac{\left[\overrightarrow{C_{2}B_{1}},\,\sum\limits_{j}\left(\vec{F}^{j}_{11}+\vec{F}^{j}_{21}\right)\right]_{z}+\left[\overrightarrow{C_{2}B_{2}},\,\sum\limits_{j}\left(\vec{F}^{j}_{22}+\vec{F}^{j}_{12}\right)\right]_{z}}{l^{2}k_{f}}\right|_{i}.$ (9)
@@
where $\omega_{z}|_{i}$ is calculated using the formula (9). The rotation is performed counterclockwise at $\Delta\alpha>0$, and counterclockwise otherwise.
Now consider the rotational motion relative to the point of contact with another particle or the fixing radius, $R_{f}$ (EIII). If the nanotube $B_{1}\,B_{2}$ at some point in time touches with its edge $B_{2}$ the fixing radius or another particle, then its further motion is only rotational motion around the point of contact. In this case, the angular velocity of rotation at each time step is determined by the formula
@@
$\left.\omega_{z}\right|_{i}=2\left.\frac{\left.\left[\overrightarrow{NB_{1}},\,\sum\limits_{j}\left(\vec{F}^{j}_{11}+\vec{F}^{j}_{21}\right)\right]\right|_{z}+\left.\left[\overrightarrow{NB_{2}},\,\sum\limits_{j}\left(\vec{F}^{j}_{12}+\vec{F}^{j}_{22}\right)\right]\right|_{z}}{l^{2}k_{f}}\right|_{i}.$ (11)
@@ II.3 Calculation algorithm
As variable parameters in this paper, we consider next of them: the number of particles, $N_{p}$, the droplet evaporation time, $t_{\mathrm{max}}$, and the charge of nanotubes (parameter $kq^{2}$, where $k$ is the constant in Coulomb’s Law, and $q$ is the charge on the nanotube). These parameters are input parameters for the program. With $N_{p}$ we can vary the initial concentration of the solution. The charge of nanotubes determines the strength of their electrostatic interaction. The time $t_{\mathrm{max}}$ affects the capillary flow velocity and the advective mass transport. The disadvantage of the analytical formula (3) for the flow velocity is the singularity at $r\to R$ and $t\to t_{\mathrm{max}}$. During the creation of the program, we take this feature into account and use a modified formula,
$v_{r}(r,t)=\begin{cases}v_{r}(r,t),&(r\leqslant 0.85R)\wedge(t\leqslant 0.98t_{\mathrm{max}});\\ v_{r}(0.85R,0.98t_{\mathrm{max}}),&(r>0.85R)\wedge(t>0.98t_{\mathrm{max}});\\ v_{r}(0.85R,t),&(r>0.85R)\wedge\neg(t>0.98t_{\mathrm{max}});\\ v_{r}(r,0.98t_{\mathrm{max}}),&\neg(r>0.85R)\wedge(t>0.98t_{\mathrm{max}}).\\ \end{cases}$
@@
The calculation of the total area of the droplet takes too long due to the large number of particles, $N_{p}$ (several months). Perhaps using parallel algorithms and computations on a supercomputer would help solve this problem. Given the axial symmetry of the problem, we can limit ourselves to a specific sector of the circle with periodic boundary conditions. In the calculations, we use the sector with angle of $\pi/128$. To simplify the implementation of the calculation of the electrostatic interaction force between particles near the boundaries of the sector, we use two adjacent sectors with nanotubes duplicating the main particles in the central sector.
@@
The smaller the value of the time step $\Delta t$ used in the calculation, the more accurate the results obtained. However, at the same time, the longer the execution time of the program will be. Taking into account both factors (accuracy of calculations and execution time), we chose the optimum value of $\Delta t=10^{-4}$ s for us. With this value, our calculations take less than two weeks using Intel(R) Xeon(R) CPU E5-2690 v3 2.6 GHz. We have compared some results for special cases with similar results obtained using $\Delta t=10^{-5}$ s and have not noticed any qualitative differences. At $\Delta t=10^{-5}$ s, the calculation takes about ten times longer. The calculation time is also influenced by other parameters, such as $N_{p}$, $t_{\mathrm{max}}$ and $kq^{2}$. The pseudocode of the algorithm is presented in the listing 1.
Due to the impossibility of the fulfilling condition of $\Delta t\to 0$ in the numerical simulations, large angles of rotation of a nanotube can occur due to a strong electrostatic impulse caused by a very close position of the current particle relative to its neighbour. This problem is caused by the discreteness of time. For each small time step, the rotation of the particle must be negligible, as it corresponds to a smooth motion without sudden jumps. This also applies to the flow rotation. For this reason, we check the condition at each time step. If $\alpha>\alpha_{\mathrm{max}}$, we assign a value $\alpha$ to $\alpha_{\mathrm{max}}$. For calculations, we chose $\alpha_{\mathrm{max}}=1^{\circ}$, having previously compared it to the case when $\alpha_{\mathrm{max}}=10^{\circ}$ (no qualitative differences have been noticed). In addition, the numerical evaluation of the rotation angle of the nanotube in the case where the particle is subjected to the maximum torque from the liquid (the nanotube is fixed by its edge at the boundary $R_{f}$) using formula (5) has shown that the maximum rotation angle is less than $1^{\circ}$.
Let each nanotube be in one of three states, with a specific color used for each particle depending on its status: movable (green), frontier (black), and motionless (red). At the initial moment, particles are placed randomly in the area to ensure that they do not overlap (without intersections). By default, all nanotubes are colored green, indicating that they are subject to various types of transfer: advection, diffusion, and electrostatic interactions. Those particles that touch the boundary of $R_{f}$ are colored black. Such particles are subject only to advective and electrostatic forces, since partial contact with the substrate is assumed. Nanotubes beyond the fixing radius ($r>R_{f}$) are colored red. This is how we mark those particles that have come into contact with the substrate (deposit).
@@
7:         Calculate the electrostatic forces for the particle $p[n]$ (see formula (8)).
@@
15:              Override the color of the particle with regard to its position relatively to $R_{f}$.
@@
20:                           Calculate new coordinates of the particle resulting from the diffusion (see formula (2)).
@@
22:                       Calculate new coordinates of the particle resulting from the advection (see formulas (3), (4)).
23:                       Calculate new coordinates of the particle resulting from the electrostatics and drag forces (see formulas (7), (9)).
@@
29:                           Calculate new coordinates of particle by rotating it against collision point (see formulas (5), (6), (10), and (11)).
@@
37:                       Calculate new coordinates of the particle resulting from rotation against its intersection point with $R_{f}$ (see formulas (5), (11)).
@@ III Results and discussion
Here, to evaluate the ordering of structures, we use the nematic order parameter $S=\left\langle\cos(2\Theta)\right\rangle$, where $\Theta$ is the angle between the nanotube and the nematic direction [23, 16]. In our case (circular domain), it is the angle between the nanotube and the perpendicular to the radius drawn from the centre of the region to the centre of the particle [23] (for details, see the link https://isanm.space). A value of $S=1$ indicates that the nanotubes are aligned parallel to the external circular boundary. At $S=-1$, they are aligned perpendicular to this boundary. The value of $S=0$ indicates a chaotic structure. Each computational experiment has been repeated ten times and then the results have been averaged. Furthermore, the error has been assessed using the standard error of the mean. By default, all three factors (advection, diffusion, and electrostatics) are included in the calculations, unless otherwise stated.
According to formula (3), $v_{r}\sim 1/t_{\mathrm{max}}$, meaning that the smaller the droplet evaporation time, the greater the capillary flow velocity, also known as compensatory flow or evaporation-induced flow. A high flow velocity results in a denser packing of nanotubes, as particles near the periphery are pressed to each other more strongly by the flow compared to a slow flow. (Fig. 3). Electrostatic repulsion of nanotubes counteracts this compression more noticeably in the case of a slow flow. Increasing the parameter $t_{\mathrm{max}}$ allows one to obtain a wider deposit layer near the periphery. When transitioning from considering a designated sector to the entire region, the focus shifts to a ring-shaped deposit. Thus, it becomes possible to control the thickness of the nanotube ring. With a relatively strong particle charge ($kq^{2}=10^{-26}$ m2N) and the time of $t_{\mathrm{max}}=$ 200 s, we observe a transition from a chaotic region to an ordered one (moving from the center to the periphery), similar to the experiment [8]. In the study  [8], the nanotubes had a negative charge, but this sign is omitted here as it does not affect the result.
@@
Without taking into account diffusion, the model predicts the chaotic structure (Fig. 4). There is no transition to order in this case. This indicates that diffusion contributes by allowing the nanotubes to arrange more densely. With a relatively low charge ($kq^{2}=10^{-28}$ m2N) many local voids are observed in the structure (Fig. 4b). The shape of some of them is close to that of the circle, with a diameter several times the length of the nanotube. Such voids were not observed in the experiment [8]. The fact is that a strong charge allows nanotubes to align with each other beforehand. As a result, such noticeable voids are not formed (Fig. 4a).
@@
Figure 4: Calculation results for $N_{p}=3000$ and $t_{\mathrm{max}}=$ 200 s considering only advection and electrostatics: (a, b) deposit structure and (c, d) order parameter $S$ for $kq^{2}=10^{-26}$ m2N (a, c) and $kq^{2}=10^{-28}$ m2N (b, d).
@@
With a small charge ($kq^{2}=10^{-28}$ m2N), the structure turns out to be more dense (Fig. 5), than with a large one ($kq^{2}=10^{-26}$ m2N), see (Fig. 3). It means that the morphology of the deposit is also determined by the confrontation between advection and electrostatics, as it has been already noted above in the text. With decreasing $t_{\mathrm{max}}$ or $kq^{2}$, advection begins to dominate electrostatics. In the case of an increase in the values of these parameters, the influence of electrostatics becomes more noticeable. As a result, advection fades into the background. For this reason, the width of the nanotube layer becomes thinner (Fig. 5). Note that nonlinear behavior of $S$ is observed at $t_{\mathrm{max}}=$ 2000 s (Fig. 5c, 5f). Such layered morphology qualitatively resembles the results of other experimental work (see Fig. 6 [4]).
@@
Figure 5: The calculation results for $N_{p}=3000$ taking into account advection, diffusion, and electrostatics ($kq^{2}=10^{-28}$ m2N): (a)–(c) deposit structure and (d)–(f) order parameter $S$ for $t_{\mathrm{max}}=$ 20 s (a, d), 200 s (b, e) and 2000 s (c, f).
When considering diffusion, with increasing concentration and charge, the formation of local ordered stack structures (Fig. 6) becomes more pronounced, gradually arising, moving and disappearing (see Supplementary Material B). Similar structures were previously observed in a lattice model describing self-organization in a 2D system of rods with two mutually perpendicular orientations using random walk [24]. The fixing radius is indicated by a blue line in Fig. 6.
@@
Figure 6: Formation of stack structures at different points of evaporation time: (a) $t=$ 145 s and $t=$ 190 s (calculation parameters: $N_{p}=4000$, $kq^{2}=10^{-26}$ m2N, $t_{\mathrm{max}}=200$ s).
@@
There are no qualitative differences in the order parameter, $S$, are observed for different values $N_{p}$ (see Supplementary Material A). In the experiment [8], it was shown that, at a low concentration of the solution, alignment of the nanotubes was not observed. The proposed 2D model does not allow this observation to be reproduced. To do this, it is necessary to further develop a 3D model with a multi-layer deposit. The 2D model presented here predicts that the width of the deposit, $w$, grows with increasing initial concentration of the solution. There is a parabolic dependence of $w$ on $N_{p}$ (Fig. 7). It should be noted that, in the results presented here, $x$, $y$, and $w$ are scaled by $R$.

### v2

@@ II.2.4 Electrostatic interaction and viscous friction
Within the scope of this proposed model, the motion of a particle in the flow is superimposed onto the motion of the particle relative to the liquid under the influence of both electrostatic and viscous forces. Consider separately the possible types of motion of the nanotube under the influence of electrostatic forces: (EI) translational motion, (EII) free rotational motion relative to the center of mass of the particle, and (EIII) rotational motion relative to a point of contact with another particle or the fixing radius, $R_{f}$.
Suppose that at each small time step, the nanotube under consideration undergoes translational motion (EI) with a constant velocity $\vec{u}^{\mathrm{rod}}$ (steady motion is the result of the drag force balancing the electrostatic force):
$\vec{F}-k_{f}\vec{u}^{\mathrm{rod}}=0,$
$\Rightarrow\quad\left.u^{\mathrm{rod}}_{x}\right|_{i}=\left.\left(\sum\limits_{j}F_{x}^{j}\right)/k_{f}\right|_{i},\,\left.u^{\mathrm{rod}}_{y}\right|_{i}=\left.\left(\sum\limits_{j}F_{y}^{j}\right)/k_{f}\right|_{i}.$ (7)
Here, $\vec{F}^{j}$ is the electrostatic force acting on the current nanotube from the $j$-th particle. For each particle, we calculate the resultant force by summing interactions with all adjacent nanotubes. The viscous friction force is considered to be proportional to the first power of the particle velocity relative to the medium. The viscous friction coefficient $k_{f}$ is described in detail in Appendix B. To simplify calculations, we assume that the charge of a nanotube is concentrated at its ends (two identical point charges, $q$). The exact formula for calculating the electrostatic force in the case when the charge is evenly distributed over the nanotube can be found in Supplemental Material A (Sec. 1). However, implementing this formulation presents significant computational challenges in practice.
@@
The force exerted by nanotube $I$ ($A_{1}\,A_{2}$) on nanotube $II$ ($B_{1}\,B_{2}$) is equal to the sum of four forces, each of which is determined by Coulomb’s law
@@
For example, $\vec{F}_{11}=kq^{2}\frac{\overrightarrow{A_{1}B_{1}}}{|A_{1}B_{1}|^{3}}$, where $k$ is the constant in Coulomb’s law. The rest of the forces are recorded in the same way. During the time interval $\Delta t$, the particle moves by $\left.u^{\mathrm{rod}}_{x}\right|_{i}\times\Delta t$ along the $OX$ axis and by $\left.u^{\mathrm{rod}}_{y}\right|_{i}\times\Delta t$ along the $OY$ axis.
Consider the free rotational motion relative to the center of mass of the nanotube (EII). Let us write the moment of forces $\vec{M}$, acting on particle $II$ ($B_{1}\,B_{2}$) from the neighboring particle $I$ ($A_{1}\,A_{2}$), relative to the center of the nanotube corresponding to point $C_{2}$ (see Fig. 2),
@@
Let us assume that at each small time step $i$ the moment of resistance balances the moment of electrostatic force, $\sum_{j}\vec{M}^{j}=\vec{M}_{\mathrm{vis}}|_{i}$. Then, the particle $B_{1}\,B_{2}$ rotates around the center of mass at a constant angular velocity at each time step:
$\left.\omega_{z}\right|_{i}=\left.\frac{\left[\overrightarrow{C_{2}B_{1}},\,\sum\limits_{j}\left(\vec{F}^{j}_{11}+\vec{F}^{j}_{21}\right)\right]_{z}+\left[\overrightarrow{C_{2}B_{2}},\,\sum\limits_{j}\left(\vec{F}^{j}_{22}+\vec{F}^{j}_{12}\right)\right]_{z}}{l^{2}k_{f}/4}\right|_{i}.$ (9)
@@
where $\omega_{z}|_{i}$ is calculated using the formula (9). The rotation is performed counterclockwise at $\Delta\alpha>0$ and clockwise otherwise.
Now consider the rotational motion relative to the point of contact with another particle or the fixing radius, $R_{f}$ (EIII). If the nanotube $B_{1}\,B_{2}$ at some point in time contacts with its edge $B_{2}$ the fixing radius or another particle, then its further motion is only rotational motion around the point of contact. In this case, the angular velocity of rotation at each time step is determined by the formula
@@
$\left.\omega_{z}\right|_{i}=\left.\frac{\left.\left[\overrightarrow{NB_{1}},\,\sum\limits_{j}\left(\vec{F}^{j}_{11}+\vec{F}^{j}_{21}\right)\right]\right|_{z}+\left.\left[\overrightarrow{NB_{2}},\,\sum\limits_{j}\left(\vec{F}^{j}_{12}+\vec{F}^{j}_{22}\right)\right]\right|_{z}}{l^{2}k_{f}/2}\right|_{i}.$ (11)
@@ II.3 Calculation algorithm
As variable parameters in this paper, we consider next of them: the number of particles, $N_{p}$, the droplet evaporation time, $t_{\mathrm{max}}$, and the charge of nanotubes (parameter $kq^{2}$, where $k$ is the constant in Coulomb’s Law, and $q$ is the charge on the nanotube). These parameters serve as user-defined inputs to the simulation program. With $N_{p}$ we can vary the initial concentration of the solution. The charge of nanotubes determines the strength of their electrostatic interaction. The time $t_{\mathrm{max}}$ affects the capillary flow velocity and the advective mass transport. The disadvantage of the analytical formula (3) for the flow velocity is the singularity at $r\to R$ and $t\to t_{\mathrm{max}}$. During the creation of the program, we take this feature into account and use a modified formula
$v_{r}(r,t)=\begin{cases}v_{r}(r,t),&(r\leqslant r_{\mathrm{cut}})\wedge(t\leqslant t_{\mathrm{cut}});\\ v_{r}(r_{\mathrm{cut}},t_{\mathrm{cut}}),&(r>r_{\mathrm{cut}})\wedge(t>t_{\mathrm{cut}});\\ v_{r}(r_{\mathrm{cut}},t),&(r>r_{\mathrm{cut}})\wedge\neg(t>t_{\mathrm{cut}});\\ v_{r}(r,t_{\mathrm{cut}}),&\neg(r>r_{\mathrm{cut}})\wedge(t>t_{\mathrm{cut}}).\\ \end{cases}$ (12)
In Formula (12), we use the values of the parameters $r_{\mathrm{cut}}$ and $t_{\mathrm{cut}}$, so that the maximum value of flow velocity $v_{\mathrm{max}}$ is of the same order of magnitude as the value obtained from the experiment with a water droplet located on a substrate [16] ($R\approx$ 0.45 mm, $t_{\mathrm{max}}\approx$ 110 s). In Ref. [16], the measurements showed $v_{\mathrm{max}}\approx$ 60 $\mu$m/s, while Formula (12) outputs the value $v_{\mathrm{max}}\approx$ 97.5 $\mu$m/s at the corresponding values of $R$ and $t_{\mathrm{max}}$ for the parameters $r_{\mathrm{cut}}=0.85R$ and $t_{\mathrm{cut}}=0.98t_{\mathrm{max}}$ used here in the calculations. The effect of the parameters $r_{\mathrm{cut}}$ and $t_{\mathrm{cut}}$ on the result is discussed in Supplemental Material A (see comments to Fig. 6 in Sec. 3).
@@
The calculation of the total area of the droplet takes too long due to the large number of particles, $N_{p}$ (several months). Perhaps using parallel algorithms and computations on a supercomputer would help solve this problem. Given the axial symmetry of the problem, we can limit ourselves to a specific sector of the circle with periodic boundary conditions. In the calculations, we use the sector with the angle of $\delta=\pi/128$ (the effect of $\delta$ on the simulation results is discussed in Supplemental Material A, Sec. 3). To simplify the implementation of the calculation of the electrostatic interaction force between particles near the boundaries of the sector, we use two adjacent sectors with nanotubes duplicating the main particles in the central sector.
@@
The smaller the value of the time step $\Delta t$ used in the calculation, the more accurate the results obtained. However, at the same time, the longer the execution time of the program will be. Taking into account both factors (accuracy of calculations and execution time), we chose the optimum value of $\Delta t=10^{-4}$ s for us. With this value, our calculations take less than two weeks using Intel(R) Xeon(R) CPU E5-2690 v3 2.6 GHz. We have compared some results for special cases with similar results obtained using $\Delta t=10^{-5}$ s and have not noticed any qualitative differences (see Supplemental Material A, Sec.  3). At $\Delta t=10^{-5}$ s, the calculation takes about 10 times longer. The calculation time is also influenced by other parameters, such as $N_{p}$, $t_{\mathrm{max}}$ and $kq^{2}$. The pseudocode of the algorithm is presented in listing 1.
@@
7:         Calculate the electrostatic forces for the particle $p[n]$ [see formula (8)].
@@
15:              Override the color of the particle with regard to its position relative to $R_{f}$.
@@
20:                           Calculate new coordinates of the particle resulting from the diffusion [see formula (2)].
@@
22:                       Calculate new coordinates of the particle resulting from the advection [see formulas (3), (4), and (12)].
23:                       Calculate new coordinates of the particle resulting from the electrostatics and drag forces [see formulas (7) and (9)].
@@
29:                           Calculate new coordinates of particle by rotating it against collision point [see formulas (5), (6), (10), and (11)].
@@
37:                       Calculate new coordinates of the particle resulting from rotation against its intersection point with $R_{f}$ [see formulas (5) and (11)].
@@
Due to the impossibility of fulfilling the condition of $\Delta t\to 0$ in the numerical simulations, large angles of rotation of a nanotube can occur due to a strong electrostatic impulse caused by a very close position of the current particle relative to its neighbor. This problem is caused by the discreteness of time. For each small time step, the rotation of the particle must be negligible, as it corresponds to a smooth motion without sudden jumps. This also applies to the flow rotation. For this reason, we check the condition at each time step. If $\alpha>\alpha_{\mathrm{max}}$, we assign a value $\alpha$ to $\alpha_{\mathrm{max}}$. For calculations, we chose $\alpha_{\mathrm{max}}=1^{\circ}$, having previously compared it to the case when $\alpha_{\mathrm{max}}=10^{\circ}$ (no qualitative differences have been noticed). In addition, the numerical evaluation of the rotation angle of the nanotube in the case where the particle is subjected to the maximum torque from the liquid (the nanotube is fixed by its edge at the boundary $R_{f}$) using formula (5) has shown that the maximum rotation angle is less than $1^{\circ}$. Additional discussion of $\alpha_{\mathrm{max}}$ usage is provided in Supplemental Material A, Sec.  3.
Let each nanotube be in one of three states, with a specific color used for each particle depending on its status: movable (green), frontier (black), and motionless (red). At the initial moment, particles are placed randomly in the area to ensure that they do not overlap (without intersections). By default, all nanotubes are colored green, indicating that they are subject to various types of transfer: advection, diffusion, and electrostatic interactions. Those particles that touch the boundary of $R_{f}$ are colored black. Such particles are subject only to advective and electrostatic forces, since partial contact with the substrate is assumed. Nanotubes beyond the fixing radius ($r>R_{f}$) are colored red. This color label identifies particles that have established contact with the substrate (deposit).
We consider only pairwise interactions between the target particle and its neighbors within a cutoff radius $R_{\mathrm{cut}}=3l$. In reality, such interaction is long-range. But at a distance of $3l$, the interaction force between two point charges is almost an order of magnitude less than at a distance of $l$ (the effect of $R_{\mathrm{cut}}$ on the simulation results is discussed in Supplemental Material A, Sec.  3).
@@ III Results and discussion
Here, to evaluate the ordering of structures, we use the nematic order parameter $S=\left\langle\cos(2\Theta)\right\rangle$, where $\Theta$ is the angle between the nanotube and the nematic direction [25, 18]. In our case (circular domain), it is the angle between the nanotube and the perpendicular to the radius drawn from the center of the region to the center of the particle [25] (for details, see the link https://isanm.space). A value of $S=1$ indicates that the nanotubes are aligned parallel to the external circular boundary. At $S=-1$, particles are aligned perpendicular to this boundary. The value of $S=0$ indicates a chaotic structure. Each computational experiment has been repeated 10 times and then the results have been averaged. Furthermore, the error has been assessed using the standard error of the mean. By default, all three factors (advection, diffusion, and electrostatics) are accounted for in the calculations, unless otherwise stated.
According to formula (3), $v_{r}\sim R/t_{\mathrm{max}}$, meaning that the smaller the droplet evaporation time, the greater the capillary flow velocity, also known as compensatory flow or evaporation-induced flow. A high flow velocity results in a denser packing of nanotubes, as particles near the periphery are pressed to each other more strongly by the flow compared to a slow flow (Fig. 3). Electrostatic repulsion of nanotubes counteracts this compression more noticeably in the case of a slow flow. Increasing the parameter $t_{\mathrm{max}}$ allows one to obtain a wider deposit layer near the periphery. When the analysis is extended from a specified sector to the entire domain, the deposition pattern resolves into a ring-shaped one. Thus, it becomes possible to control the thickness of the nanotube ring. With a relatively strong particle charge ($kq^{2}=10^{-26}$ m2N) and the time of $t_{\mathrm{max}}=$ 200 s, we observe a transition from a chaotic region to an ordered one (moving from the center to the periphery), similar to the experiment (see Fig. 3 in Ref. [8]). In the study [8], the nanotubes had a negative charge, but this sign is omitted here as it does not affect the result.
@@
Without taking into account diffusion, the model predicts the chaotic structure (Fig. 4). There is no transition to order in this case. This indicates that diffusion contributes by allowing the nanotubes to arrange more densely. With a relatively low charge ($kq^{2}=10^{-28}$ m2N) many local voids are observed in the structure [Fig. 4(b)]. The shape of some of them is close to that of a circle, with a diameter several times the length of the nanotube. Such voids were not observed in the experiment [8]. The strong electrostatic interaction induces prealignment of nanotubes prior to deposition. As a result, such noticeable voids are not formed [Fig. 4(a)].
@@
Figure 4: Calculation results for $N_{p}=3000$ and $t_{\mathrm{max}}=$ 200 s considering only advection and electrostatics: (a),(b) deposit structure and (c),(d) order parameter $S$ for $kq^{2}=10^{-26}$ m2N (a),(c) and $kq^{2}=10^{-28}$ m2N (b),(d).
@@
With a small charge ($kq^{2}=10^{-28}$ m2N), the structure turns out to be more dense (Fig. 5), than with a large one ($kq^{2}=10^{-26}$ m2N), see Fig. 3. It means that the morphology of the deposit is also determined by the confrontation between advection and electrostatics, as it has been already noted above in the text. With decreasing $t_{\mathrm{max}}$ or $kq^{2}$, advection begins to dominate electrostatics. In the case of an increase in the values of these parameters, the influence of electrostatics becomes more noticeable. As a result, advection fades into the background. For this reason, the width of the nanotube layer becomes thinner (Fig. 5). Note that nonlinear behavior of $S$ is observed at $t_{\mathrm{max}}=$ 2000 s [Figs. 5(c) and 5(f)]. Such layered morphology qualitatively resembles the results of other experimental work (see Fig. 6 [4]).
@@
Figure 5: The calculation results for $N_{p}=3000$ taking into account advection, diffusion, and electrostatics ($kq^{2}=10^{-28}$ m2N): (a)–(c) deposit structure and (d)–(f) order parameter $S$ for $t_{\mathrm{max}}=$ 20 s (a),(d), 200 s (b),(e) and 2000 s (c),(f).
Refer to caption(b)
@@
Figure 6: Formation of stack structures at different points of evaporation time: (a) $t=$ 145 s and (b) $t=$ 190 s (calculation parameters: $N_{p}=4000$, $kq^{2}=10^{-26}$ m2N, $t_{\mathrm{max}}=200$ s).
When considering diffusion, with increasing concentration and charge, the formation of local ordered stack structures (Fig. 6) becomes more pronounced, gradually arising, moving and disappearing (see Supplemental Material B). Similar structures were previously observed in a lattice model describing self-organization in a 2D system of rods with two mutually perpendicular orientations using random walk [26]. The fixing radius is indicated by a blue line in Fig. 6.
@@
There are no qualitative differences in the order parameter, $S$, observed for different values $N_{p}$ (see Supplemental Material A, Sec. 2). In the experiment [8], it was shown that, at a low concentration of the solution, alignment of the nanotubes was not observed. The proposed 2D model does not allow this observation to be reproduced. To do this, it is necessary to further develop a 3D model with a multilayer deposit. The 2D model presented here predicts that the width of the deposit, $w$, grows with increasing initial concentration of the solution. There is a parabolic dependence of $w$ on $N_{p}$ (Fig. 7). It should be noted that, in the results presented here, $x$, $y$, and $w$ are scaled by $R$.